XYZ Co. operates in a competitive market. Its Total Product (Q) is given as Q=f(L)= 2L, and it takes the wage and price as given. Derive the firm's short-run demand for labor as a function of w and p. How much labor will the firm hire if W=₵20 and P=₵100?


1
Expert's answer
2021-11-04T13:40:36-0400

The short-run demand curve is given by "MP_L=w\/p=w" which is obtained by dividing the nominal demand curve by the product price, p.

"MRP_L=w"

"MP_L\times P=w"

"MP_L=2L"

"P=100"

"W=20"

"2L\times 100=20"

"200L=20"

"L=0.1"

"Q=2L=2\times 0.1=0.2"

Labor"=0.2"